Operations Report – November 05th

Crew 214 Operations Report 5-11-2019
SOL: 9
Name of person filing report: Andrew Wheeler
Non-nominal systems: Opportunity, Suit 10
Notes on non-nominal systems: Opportunity brakes do not work, Suit 10 is not functioning
Generator: run
Hours run: 11h 15m
From what time last night: 1840
To what time this morning: 0655
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before generator is run at night): 65%
Diesel Reading – 60%
Station Propane Reading – 81%
Ethanol Free Gasoline: N/A
Water (loft tank): 47 gallons
Water Meter: 1455759 units
Water (static tank): 223 gallons
Static to Loft Pump used – yes
Water in Green Hab: 225 gallons
Water in Science Dome: 0 gallons
Toilet tank emptied: no
Deimos rover used: not at hab
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –
Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –
Spirit rover used: yes
Hours: 108.0
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): 58%
Currently charging: yes
Opportunity rover used: no
Hours: 65.5
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes
Curiosity rover used: yes
Hours: 113.9
Beginning charge: (Before EVA) : 100%
Ending charge: (On return from EVA, before charging): 56%
Currently charging: yes
Notes on rovers: no
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
ATV Fuel Used: 0 Gals
# Hours the ATVs were used today: 0
Notes on ATVs: Assigned to the director
HabCar used and why, where? N/A
CrewCar used and why, where? N/A
General notes and comments: N/A
Summary of internet: Functional
Summary of suits and radios: 5 Radios and 3 earpieces that need servicing have been identified and set aside, the battery for suit 10 is fully charged and it is Suit 10 that is not functional, further information on Suit 5,s helmet has been forwarded to Capcom.
Summary of Hab operations: All systems nominal
Summary of GreenHab operations: Atmospheric equalization to the green hab was carried out during the day, plants were watered..
Summary of ScienceDome operations: All systems nominal. Sample processing following EVA was carried out.
Summary of RAM operations: Not used
Summary of any observatory issues: N/A
Summary of health and safety issues: All crew healthy
Questions, concerns and requests to Mission Support: Between 20191104:1140 and 20191105: 1100, the micrometeorite grid has been visited and trampled by at least one smallish adult (wearing boots). This has compromised a number of the plots in the unsampled area and will affect the integrity of the experiment. Whilst this is disappointing, we realize we are on public land and, in retrospect, we should have erected a temporary sign and barrier stating “please keep clear – monitoring in progress” to prevent this. We believe future rotations may benefit from this advice. We note that our daily reports identify the specific location where our field research is undertaken.

Copyright © The Mars Society. All rights reserved. | Main Site